lolokrat:

can someone explain how to find these without using a calculator? my prof gave us answers to them but no explanations

lolokrat:

Extrinix:

I'll do the first one as an example \(log_3 27=x\) Definition of logarithmic \(3^x=27\) Find exponents with equal exponentials \(3^{1x}=3^3\) Cancel the exponentials, leaving you with powers \(1x=3\) Solve for x \(x=\dfrac{3}{1}\) \(x=3\)

lolokrat:

thank you so much!